OpenStudy (anonymous):

find the real number of solutions of the equation x^3-6x=0

OpenStudy (anonymous):

x^3-6x=0 x(x^2-6)=0 Either x=0 or x^2-6= x^2=6 So \[x=\pm \sqrt{6}\]

OpenStudy (anonymous):

okay thanks so much!